OpenStudy (anonymous):

solve the equation for x. radical 4x+57=x+3

OpenStudy (anonymous):

4x + 57 = x + 3 4x + x = 57 + 3 5x = 60 5x/5 = 60/5 x = 12

OpenStudy (anonymous):

\[\sqrt{4x + 57} = x + 3\] Start with squaring both sides to get:\[4x + 57 = (x + 3)^2\] This gets us: \[4x + 57 = (x+3)(x+3)\] which is:\[4x + 57 = x^2 + 6x + 9\] Moving everything to the right side we get:\[0 = x^2 - 2x - 48\] This factors into \[0 = (x-8)(x-6)\] Setting both of those equal to 0 gets you x = 8 and x = 6

OpenStudy (anonymous):

I think you got a sign wrong there: should be:\[x ^{2}+2x-48\] which factors to (x+8)(x-6)